Database last updated on Sun, 01 Feb 2026 09:40:50 +0000 by localhost
Instrument constant changes for the week before 2024011 (2 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 7 rows)
HG burned out? (AP results when HG is ON) (0 sec, 10 rows)
Measurements and tests summary.
| _jday | _loc | _serial | _skc | dt | ed | hg | rs | sl | zs |
| 011 | TORONTO | #017 | cal,Next schedule will be cal (now cal ZS= 85 ZS= 85.516),The next day will run schedule cal,Waiting until end of local day.,Waiting until ZA = | 4 | 2 | 10 | 4 | 4 | 91 |
(1 sec)
Number of days since last dome/window cleaning (0 sec, 2 rows)
| serial | days |
| 017 | more than a year |
Missing files (0 sec, 1 rows)
| Serial | last data | days ago |
| 017 | 2024012 | 0-1 |
Questionable Run/Stop results. (0 sec, 1 rows)
| Serial | Jday | slit 0 | dark rs | dark counts | slit 1 | slit 2 | slit 3 | slit 4 | slit 5 | slits 3+5 |
| 017 | 2024011 | 0.997 | 5.000 | 2 / 0 (1.5M) | 0.996 | 0.994 | 0.999 | 0.999 | 1.000 | 0.999 |
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 017 | 28 | 2 | 12 | 4 | 1 | 0 | 2 | 1 | 0 |
AP results for MB Brewer. (0 sec, 1 rows)
| Serial | Jday | t1 off | t1 on | t2 off | t2 on | t3 off | t3 on | HV off | HV on | +15V off | +15V on | +5V off | +5V on | -15V off | -15V on | +24V off | +24V on | +5V sps off | +5V sps on | -8V sps off | -8V sps on | SL A off | SL A on | SL V off | SL V on | W |
| 017 | 2024011 | 4.5 | 4.5 | 5.8 | 5.8 | 5.2 | 5.2 | 1320.7 | 1315.2 | 15.1 | 14.9 | 5.1 | 5.1 | -15.0 | -15.0 | 24.2 | 24.3 | 4.9 | 4.9 | -8.0 | -8.0 | 9.1 | 1.6 | 0.0 | 10.36 | 16.2 |
Suspicious HG test results (correlation<0.98 or counts<5000). (0 sec, 1 rows)
| Serial | Jday | count | Peak | min.corr. | max.corr. | 30-day history |
| #017 | 011 | 1 | 37048 | 0.979 | 0.979 |